Get the mean, median, and mode of the following data below. Write only the correct answer.

Age of the 50 Respondents

20 26 27 19 17 17 18 16 21 23

22 26 25 24 25 24 21 20 21 25

27 25 19 18 16 17 25 26 26 21

22 24 24 26 27 19 18 16 15 15

19 15 20 25 26 25 27 15 24 27


1
Expert's answer
2020-12-15T13:36:42-0500

mean

"\bar x=\frac{\sum x}{n}"

="\frac{20+26+27+...+27}{50}"

="\frac{1086}{50}"

=21.72

median

Arrange the numbers from the smallest to the largest.

the 25th and the 26th values are both 22. (22+22)/2=22

=22

mode

The most appearing number is 25. it appears 7 times.

=25